Claire's cousin is now 3 times as old as Claire. In 10 years, her cousin will be twice as old as Claire. How old is Claire now?

Found 2 solutions by Alan3354, ikleyn:
Answer by Alan3354(69443)

30 and 10 now.

Answer by ikleyn(52782)

.
Claire's cousin is now 3 times as old as Claire. In 10 years, her cousin will be twice as old as Claire. How old is Claire now?
~~~~~~~~~~~~~~


Clara = x  years old

The cousin is 3x years old now.



In 10 years

    3x + 10 = 2*(x+10)



Simplify and solve

    3x + 10 = 2x + 20

    3x - 2x = 20 - 10

       x    = 10.


ANSWER.  Clara is 10 years old now.

Solved.